# Turbulent dynamos in a collapsing cloud

### v1

@@ II.2 The large-scale dynamo (LSD)
[…]
where $\tilde{\bm{\omega}}=\tilde{\nabla}\times\tilde{\mathbf{v}}$ and $\tau_{0}$ is the correlation time of the random flow. Assuming for simplicity that $\langle\mathbf{V}\rangle$ represents a solid-body rotation, the averaged induction equation written in the rotating frame follows as
$\frac{\partial\langle\tilde{\mathbf{B}}\rangle}{\partial\tilde{t}}=\tilde{\nabla}\times\left[\tilde{\alpha}\langle\tilde{\mathbf{B}}\rangle\tilde{\eta}_{\text{T}}\tilde{\nabla}\times\langle\tilde{\mathbf{B}}\rangle\right],$ (13)
@@
where $\tilde{k}$ is the wavenumber of the mean magnetic field. A similar expression is valid for a spherically symmetric distribution of $\tilde{\alpha}$ [54]. The magnetic field of the scale $2\pi/\tilde{k}_{\text{m}}=4\pi\tilde{\eta}_{\text{T}}/\tilde{\alpha}$ grows most rapidly as $\exp(\tilde{\gamma}_{\text{m}}\tilde{t})$ with $\tilde{\gamma}_{\text{m}}=\tilde{\gamma}(\tilde{k}_{\text{m}})=\tilde{\alpha}^{2}/(4\tilde{\eta}_{\text{T}})$.
@@
III The forcing function
$\tilde{\mathbf{f}}\left(\tilde{\mathbf{x}},\tilde{t}\right)$ $=\mathrm{Re}\left\{N\mathbf{f}_{\tilde{\mathbf{k}}(\tilde{t})}\;\exp{\left[\mathrm{i}\tilde{\mathbf{k}}(\tilde{t})\cdot\tilde{\mathbf{x}}+\mathrm{i}\phi(\tilde{t})\right]}\right\},$ (14)
where $\tilde{\mathbf{k}}$ is the supercomoving wave-vector. The normalization factor $N$ decreases with the time step $\Delta\tilde{t}$ to ensure that the forcing is approximately $\updelta$-correlated in the supercomoving time. The time step is calculated using the CFL condition. The remaining factors in $N$ are fixed using dimensional arguments with a dimensionless parameter $f_{0}$ to control the forcing strength; we set $f_{0}=1$. The factor $\sigma$ in $R_{ij}$ is a measure of the kinetic helicity. For SSD we use $\sigma=0$ to drive a non-helical flow, whereas $\sigma=1$ for LSD to generate a helical velocity field required to produce a large-scale magnetic field.
IV Duration of the kinematic stage
@@
with $t_{\rm ff}$ the free-fall time and $\tilde{\gamma}$ the growth rate. The time at which kinematic regime ends for the dynamo in a stationary background (the ‘standard’ dynamo) is related to $\tilde{\gamma}$ and $B_{0}$ as
@@
Figure 1: Comparison of the evolution of $B_{\rm rms}$ in the SSD for the numerical resolutions $128^{3}$ and $96^{3}$.
Figure 2: As Fig. 1 but for the LSD. The inset shows the slow transition to the dynamo saturation completing around $t/t_{0}=160$.
@@ V Numerical resolution tests
Here we verify that the numerical resolution of $128^{3}$ is sufficient for our purposes. Figures 1 and 2 show the evolution of the rms physical magnetic field strength $B_{\rm rms}$ in two simulations of SSD and LSD, respectively, with the resolutions $128^{3}$ and $96^{3}$. The simulation parameters are the same as described in the main text. We find the agreement between the results obtained under the two numerical resolutions to be sufficiently close to justify $128^{3}$ as the resolution used to obtain our main results.

### v2

@@ I.2 The large-scale dynamo (LSD)
$\tilde{\alpha}\approx-\frac{1}{3}\tilde{\tau}_{0}\left<\tilde{\mathbf{v}}\cdot\tilde{\bm{\omega}}\right>,\quad\tilde{\eta}_{t}\approx\frac{1}{3}\tilde{\tau}_{0}\left<\tilde{\mathbf{v}}^{2}\right$>, (24)
where $\tilde{\bm{\omega}}=\tilde{\nabla}\times\tilde{\mathbf{v}}$ and $\tau_{0}$ is the correlation time of the random flow. The derivation assumes a small $\widetilde{\tau}_{0}$ and $R_{\text{m}}\gg 1$, such that the resistive and nonlinear terms in the fluctuations are negligible in comparison to the time derivative term. Assuming for simplicity that $\langle\mathbf{V}\rangle$ represents a solid-body rotation, the averaged induction equation written in the rotating frame follows as
$\frac{\partial\langle\tilde{\mathbf{B}}\rangle}{\partial\tilde{t}}=\tilde{\nabla}\times\left[\tilde{\alpha}\langle\tilde{\mathbf{B}}\rangle\;{-}\;\tilde{\eta}_{\text{T}}\tilde{\nabla}\times\langle\tilde{\mathbf{B}}\rangle\right$], (25)
@@
[…]
@@
II Duration of the kinematic stage
@@
with $t_{\rm ff}$ the free-fall time and $\tilde{\gamma}$ the growth rate. The time at which kinematic regime ends for the dynamo in a stationary background (the ‘standard dynamo’) is related to $\tilde{\gamma}$ and $B_{0}$ as
@@
III Dynamo action of decaying turbulence
Super-exponential amplification of magnetic field can occur even if turbulence in a collapsing cloud is not continuously driven but rather decays. Thus, we consider the case of an initially turbulent plasma, with no forcing and a weak seed magnetic fields undergoing gravitational collapse. Now we have decaying turbulence, unlike the case of a driven turbulence explored in the main text.
The supercomoving (SC) MHD equations in the kinematic stage (when the Lorentz force is negligible) are same as the standard MHD equations. Since the initial magnetic field is weak, the turbulence decays as in a purely hydrodynamic case. In a Kolmogorov turbulence, the rate of decay of supercomoving kinetic energy is given by
$\frac{\mathrm{d}\tilde{v}^{2}}{\mathrm{d}\tilde{t}}=-{\eta_{\rm d}}\;\frac{\tilde{v}^{3}}{\tilde{l}}$, (31)
where $\eta_{\rm d}$ is the dissipation factor [34].
From the conservation of either the angular momentum (Loitsyansky invariant) or the linear momentum (Saffman invariant), we have the following constraint on the evolution of the outer turbulent scale:
$\tilde{v}_{0}^{\alpha}\tilde{l}_{0}=\text{const}={\tilde{v}_{0i}^{\alpha}\tilde{l}_{0i}=\psi_{0},}$ (32)
where $\alpha=2/5$ if the Loitsyansky invariant applies and $\alpha=2/3$ in the case of the Saffman invariant [72], and the subscript $i$ denotes the initial values. We assume a peaked velocity spectrum as is standard in decaying turbulence literature. Thus we can combine Eq. (32) with Eq. (31) to obtain $\tilde{v}_{0}\propto\tilde{t}^{-{1}/{\left(1+\alpha\right)}}$. Substituting this in Eq. (32) gives the evolution of the length scale as $\tilde{l}_{0}\propto\tilde{t}^{{\alpha}/{\left(1+\alpha\right)}}.$
Magnetic field amplification is possible if the system is super-critical with respect to the dynamo action, that is magnetic Reynolds number is larger than the critical Reynolds number, $R_{\text{m}}>R_{\text{m}}^{\text{cr}}$. The amplification continues as long as the eddies available in the spectrum remain super-critical. During this period, the magnetic field strength grows as
$a^{2}B=B_{0}\exp{\left(\int\tilde{\gamma}\,\mathrm{d}\tilde{t}\right)}$, (33)
where $B_{0}$ is the seed field strength, $a$ is the scale factor and $\tilde{\gamma}$ is the turnover rate (in the SC frame) of the fastest eddy available. As discussed in the main text, an exponential growth in the SC coordinates (which is similar to the standard kinematic dynamo) implies a super-exponential amplification in the physical frame.
In the decaying turbulence, the growth rate $\tilde{\gamma}_{0}$ due to the largest super-critical eddy of a scale $\tilde{l}_{0}$ with $R_{\text{m}}(\tilde{l}_{0})>R_{\text{m}}^{\text{cr}}$ is given by its turnover rate,
$\tilde{\gamma}_{0}=\frac{\tilde{v}_{0}}{\tilde{l}_{0}}\propto\frac{1}{\tilde{t}}$, (34)
and using the Kolmogorov scaling $\tilde{v}(\tilde{l})\propto\tilde{l}^{1/3}$ we obtain the growth rate due to eddy of scale $\tilde{l}$
$\tilde{\gamma}(\tilde{l})=\frac{\tilde{v}}{\tilde{l}}=\frac{1}{\tilde{l}}\frac{\tilde{v}_{0}\tilde{l}^{1/3}}{\tilde{l}_{0}^{1/3}}=\frac{\tilde{v}_{0}}{\tilde{l}_{0}}\left(\frac{\tilde{l}_{0}}{\tilde{l}}\right)^{2/3}.$ (35)
Focusing on the stretching of field lines due to the eddy of the size $\tilde{l}$, we fix $\tilde{l}$ to obtain the growth rate as
$\tilde{\gamma}(\tilde{l})=\frac{\tilde{v}_{0}}{\tilde{l}_{0}^{1/3}}\frac{1}{\tilde{l}^{2/3}}=\tilde{\gamma}_{C}(\tilde{l})\,\tilde{t}^{-(3+\alpha)/[3(1+\alpha)]}.$ (36)
[…]
For the case of the Loitsyansky invariant, we obtain $\alpha=2/5$, hence
$\tilde{\gamma}(\tilde{l})=\tilde{\gamma}_{C}(\tilde{l})\tilde{t}^{-{17}/{21}}.$ (37)
Equation (33) then shows that the magnetic field strength increases as
$a^{2}B$ $=B_{0}\exp{\left(\tilde{\gamma}_{C}(\tilde{l})\int\tilde{t}^{-17/21}\mathrm{d}\tilde{t}\right)}$,
$=B_{0}\exp{\left(\frac{21}{4}\tilde{\gamma}_{C}\tilde{t}^{4/21}\right)}.$ (38)
Similarly, for the Saffman-type decay we obtain $\tilde{\gamma}(\tilde{l})=\tilde{\gamma}_{C}(\tilde{l})\tilde{t}^{-11/15}$ and
${a^{2}B=B_{0}\exp{\left(\frac{15}{4}\tilde{\gamma}_{C}\tilde{t}^{4/15}\right)}.}$ (39)
The field strength in a freely decaying turbulence increases slower than in a steady supercomoving turbulence discussed in the main text, where the power of $\tilde{t}$ in the exponent is unity. From the definition of $\tilde{t}$ in Eq. (3) and Eq. (9) of the main text, it is clear that $\tilde{t}$ is a steeply increasing function of $t$. We conclude that a super-exponential magnetic field growth discussed in the main text occurs, albeit with a different rate of growth, even in this least favorable scenario, the case of decaying turbulence.
We emphasize that the magnetic field can only grow in a super-critical decaying turbulence. Once the turbulence becomes subcritical, the magnetic field will decay. But as far as astrophysical systems are considered, $R_{\text{m}}\gg R_{\text{m}}^{\text{cr}}$, and it can take a long time (say, $t_{R}$) to reduce $R_{\text{m}}$ to $R_{\text{m}}^{\text{cr}}$, thus providing sufficient time for the magnetic field amplification. When $t_{R}\gtrsim t_{\rm ff}$, we get a significant additional super-exponential growth due to the collapse. We also note that maintaining $R_{\text{m}}>R_{\text{m}}^{\text{cr}}$ for a sufficient time in numerical simulations and thus detecting the scalings of Eqs (III) and/or (39) can be difficult because of computational constraints on the numerical resolution.
At the end of the kinematic stage, the back-reaction from the Lorentz force saturates the magnetic field growth. The dependence of the magnetic field strength on the gas density in the saturated dynamo can be derived as follows.
The dependence of the flow speed on the scale factor is obtained by integrating Eqs (31) and (32),
$\int_{\tilde{v}_{0i}}^{\tilde{v}_{0}}\tilde{v}^{-(2+\alpha)}\mathrm{d}\tilde{v}$ $=-\frac{\eta_{\rm d}}{2\psi_{0}}\int_{0}^{\tilde{t}}\mathrm{d}\tilde{t}^{\prime}$,
$\tilde{v}_{0}^{-(1+\alpha)}$ $=\tilde{v}_{0i}^{-(1+\alpha)}+\frac{\eta_{\rm d}(1+\alpha)}{2\psi_{0}}\tilde{t}$,
$\tilde{v}_{0}$ $=\tilde{v}_{0i}\left[1+\frac{\eta_{\rm d}(1+\alpha)}{2\psi_{0}\tilde{v}_{0i}^{-(1+\alpha)}}\tilde{t}\right]^{-1/(1+\alpha)}.$ (40)
From $\psi_{0}=\tilde{v}_{0i}^{\alpha}\tilde{l}_{0i}$, we obtain $\psi_{0}\tilde{v}_{0i}^{-(1+\alpha)}=\tilde{l}_{0i}/\tilde{v}_{0i}=\tau_{L}$, the turnover time of the outer eddy. Using
$\tilde{t}=\frac{2}{s}\sqrt{\frac{1-a}{a}},\quad s=\frac{\pi}{2t_{\rm ff}},$
Eq. (III) can be written in terms of the scale factor as
$\tilde{v}_{0}$ $=\tilde{v}_{0i}\left[1+(1+\alpha)\chi\sqrt{\frac{1-a}{a}}\right]^{-1/(1+\alpha)},$ (41)
where $\chi=\eta_{\rm d}/s\tau_{L}\sim t_{\rm ff}/t_{\text{nl}}.$
The saturation stage of the dynamo is characterized by the equipartition between magnetic and kinetic energies (Eq. (10) of the main text),
$a(t)\frac{\tilde{B}^{2}}{8\pi}\simeq\frac{1}{2}\tilde{\rho}\tilde{v}_{0}^{2}.$ (42)
Assuming that Eq. (41) applies during the nonlinear dynamo stage, we obtain for $a\ll 1$, the scaling
$\tilde{v}_{0}\propto a^{1/[2(1+\alpha)]}\propto\rho^{-1/[6(1+\alpha)]},$ (43)
whereas the peculiar velocity field scales as
$v_{0}\propto a^{-(1+2\alpha)/2(1+\alpha)}\propto\rho^{(1+2\alpha)/6(1+\alpha)}.$ (44)
Finally, from Eq. (42), we obtain the scaling of the magnetic field strength as
$B\propto\sqrt{\rho v_{0}^{2}}\propto a^{-(4+5\alpha)/2(1+\alpha)}\propto\rho^{(4+5\alpha)/[6(1+\alpha)]}.$ (45)
In the limit of $\alpha\to 0$, this expression reproduces the flux-freezing exponent of 2/3, and the exponent 5/6 of the forced case is retrieved in the limit $\alpha\to\infty$ which corresponds to a constant $\tilde{v}$. The exponent is greater than 2/3 for any $\alpha>0$, irrespectively of the invariant that governs the decay. After the collapse, the magnetic field can be sustained by other processes such as convection in the case of stars, magnetic or fluid instabilities for accretion-disk flows and stellar feedback in the case of galaxies.
IV The forcing function
[…]
where $\tilde{\mathbf{k}}$ is the supercomoving wave-vector. The normalization factor $N$ decreases with the time step $\Delta\tilde{t}$ to ensure that the forcing is approximately $\updelta$-correlated in the supercomoving time. The time step is calculated using the CFL condition. Although the forcing function is $\delta$-correlated in $\tilde{t}$, the resulting supercomoving velocity is not. However, this forcing is sufficient to generate randomness in the flow, essential for exciting turbulent dynamos. The remaining factors in $N$ are fixed using dimensional arguments with a dimensionless parameter $f_{0}$ to control the forcing strength; we set $f_{0}=1$. The factor $\sigma$ in $R_{ij}$ is a measure of the kinetic helicity. For SSD we use $\sigma=0$ to drive a non-helical flow, whereas $\sigma=1$ for LSD to generate a helical velocity field required to produce a large-scale magnetic field.
@@ V Numerical resolution tests
Here we verify that the numerical resolution of $128^{3}$ is sufficient for our purposes. Figures 6 and 7 show the evolution of the rms magnetic field strength $B_{\rm rms}$ in two simulations of SSD and LSD, respectively, with the resolutions $128^{3}$ and $96^{3}$. The SSD simulation uses $k_{f}=2,\text{Re}=415$, and $\text{Pm}=1$, while the LSD simulation employs $k_{f}=4,\text{Re}=180$, and $\text{Pm}=0.1$. We find the agreement between the results obtained under the two numerical resolutions to be sufficiently close to justify $128^{3}$ as the resolution used to obtain our main results.
Figure 6: Comparison of the evolution of $B_{\rm rms}$ in the SSD for the numerical resolutions $128^{3}$ and $96^{3}$. Parameter values are the same as in Figs 1 and 2 of the main text but the results are presented here versus $t$ rather than $t/t_{0}$.
Figure 7: As Fig. 6 but for the LSD. The inset shows the slow transition to the dynamo saturation completing around t=200{t=200}.